OpenStudy (anonymous):

Determine which polynomial is a perfect square trinomial. 49x2 − 28x + 16 9a2 − 30a + 25 25b2 − 45b − 81 16x2 − 24x − 9

OpenStudy (zale101):

A perfect square trinomial is (ax^2 + bx +c) u need to know that A and C are positive values and always

OpenStudy (zale101):

So...calculate b^2 - 4ac and the polynomial will be a perfect square if ure equations equals to zero :)

OpenStudy (zale101):

\[b^2 - 4ac = 28^2 - 4*16*49 = 784 - 3136\] so A is not the answer

OpenStudy (zale101):

Did u understand of anything i said? You can do the same like i did for the first equation and if u get 0 that's the answer.

OpenStudy (zale101):

@jen1511 , do u want me to do more equation for u :) ?

OpenStudy (anonymous):

wait what im confused

OpenStudy (zale101):

I'll do B for u 9a^2 − 30a + 25 9a^2 = (3a)^2 25 = 5^2 = (3a - 5) ^2 another way b^2 - 4ac = 30^2 - 4x9x25 = 900 - 900 = 0

OpenStudy (zale101):

Does equation B seems a perfect square trinomial to u?

OpenStudy (anonymous):

I guess yeah!

OpenStudy (zale101):

Yes